Find the magnitude of two like parallel forces acting at a distance of 240 mm, whose resultant is 200 N, and its line of action is at a distance of 60 mm from one of the forces.?
1
Expert's answer
2021-10-29T02:33:14-0400

Since the forces are parallel, so the angle between two forces F1 and F2 are 0

By the formula

√(F1+F2+2F1*F2Cos0)=200

F1+F2=200------------(1)

Taking moment about resultant

F1*60-F2*(240-60)=0

60F1-180F2=0-------(2)

From equation 1 and 2

We get

F1=150KN

F2=50KN